TMS Offline
Feb 17, 2026
Feb. 12, 2026: The TMS system recently experienced a hardware issue with the GPU in the PC controlling the system. A replacement GPU is currently being sourced, but for the time TMS will be offline.
Feb 17, 2026
Feb. 12, 2026: The TMS system recently experienced a hardware issue with the GPU in the PC controlling the system. A replacement GPU is currently being sourced, but for the time TMS will be offline.
Aug 19, 2025
MODS Controller Upgrade – Instrument Availability Delayed
update: Feb 2nd
The MODS instruments underwent controller upgrades this summer, replacing the legacy head electronics and a set of aging computers. These upgrades were carried out to ensure the long-term reliability, robustness, …
Jun 28, 2025
On the night of June 26 UT when moving the 2.4″ longslit into the focal plane, the mask failed to insert into the focal plane despite the software thinking so. When the instrument was opened, the grabber was found to …
Apr 10, 2025
April 11, 2025: Problem with the oxygen regulator was found and resolved. The MOS mask milling machine has been returned to operation.
April 9, 2025: The laser for the MOS mask milling machine recently stopped working. Currently all mask milling …
Apr 03, 2025
Apr 2, 2025: Flaky switch on LUCI1 Filter Wheel #2 (broadband filters).
LUCI1’s Filter Wheel #2, where the broadband filters are located, is showing a flaky position sensing switch that affects all higher-numbered filters, which includes all but the zJspec …
Mar 09, 2025
Mar 13, 2025: All normal modes with LUCI1 are released for science.
N375 and N30 imaging, as well as G210 and G200 spectroscopic modes have all been released for normal science use.
Mar 9, 2025: LUCI1 All modes **NOT** …
Feb 14, 2025
Update Feb 24, 2025
On-sky seeing limited checkouts were completed on UTFeb24. The binodal astigmatism was measured and collimation tables updated. The SX ASM is released for all available seeing-limited science modes without restrictions.
Update Feb 14, 2025
ASM …
Jan 21, 2025
DX ASM in Cold Weather Update – Jan 21, 2025
Work was done in summer shutdown 2024 to attempt to address the issues causing the failures in cold temperatures with the DX ASM. Unfortunately, we have continued to see the …
Jan 11, 2025
Jan 09, 2025
The DX LUCI WFS experienced an issue with one of the filter wheels. This filter wheel is used for very bright targets only with an R magnitude brighter than 8th magnitude. LUCI full AO and LUCI ESM …
Sep 10, 2024
As of Sept. 9, 2024, LBTI’s camera cryostat (NIC) remains off the telescope for ongoing maintenance and upgrade work, with some major uncertainty on the observing modes that will be available for 2024B and 2025A. PIs of accepted 2024B programs …